OpenStudy (anonymous):

how to caluclate the temparature co-efficient of resistance of superconducting materials

OpenStudy (abb0t):

\[\frac{ ∆R }{ R_o }=\alpha∆T \] alpha is the temp coeff or resistance. if at some given standard temp \[\frac{ R-R_o }{ R_o } = \alpha∆T-T_o \] or \[R = R_o [1+\alpha(T-T_o)]\]